OpenStudy (ambermarie151):

Help please? Write a function that repersents this process: A:Take a number,x. B: Multiply it by4. C: Subtract 2 from the result Then find the inverse of this function. Does the inverse represent the reverse of the process above ? Explain why or why not.

OpenStudy (agent0smith):

You should be able to at least do the first part, writing the original function.

OpenStudy (austin.l):

Yeah! Lets see it! It tells you what to write in the problem!

OpenStudy (ambermarie151):

Would it be like x•4-2

OpenStudy (agent0smith):

Yes. y = 4x - 2 Now to find the inverse, solve that equation for x = ...

OpenStudy (austin.l):

You type faster @agent0smith .... :( I'll let you have this one! >:D You're in good hands OP

OpenStudy (ambermarie151):

Would I divide it by 2?

OpenStudy (agent0smith):

y = 4x - 2 If you needed to find x, what would you do first?

OpenStudy (ambermarie151):

Would you add 4 to both sides

OpenStudy (agent0smith):

No... you'd add 2 to both sides.

OpenStudy (ambermarie151):

So it would be y=6x

OpenStudy (austin.l):

Negative GhostRider. When you add a number without a variable you don't touch ones with variables. So, the \[4x\] would be left alone. as would the \[y\] Try again!

OpenStudy (ambermarie151):

This makes no sense dude I'm literally so confused

OpenStudy (austin.l):

(+2) y = 4x + 2 (-2) y + 2 = 4x Take it from there?

OpenStudy (austin.l):

Oh my, i made a typo. Don't listen to me! D: I quit!

OpenStudy (skullpatrol):

Let's start at what you said at the beginning. @ambermarie151 "Would it be like x•4-2" What you did was take x and 1. Multiplied by 4 2. Then you subtracted 2 To get back to x you apply the "inverse" operations in the opposite order. 1. Add 2 2. Divide by 4 Let y = 4x - 2 y + 2 = 4x - 2 + 2 y + 2 = 4x (y + 2)/4 = (4x)/4 y/4 + 1/2 = x Note: inverse operations are operations that "undo" each other, for example multiplication and division.